# Visualizing Thought: Conceptual Diagrams Enable Robust Planning in LMMs

### v1

@@ Abstract
Human reasoning relies on constructing and manipulating mental models—simplified internal representations of situations that we use to understand and solve problems. Conceptual diagrams (e.g., a sketch drawn by a human to aid reasoning) externalize these mental models, abstracting irrelevant details to efficiently capture relational and spatial information. In contrast, Large Language Models (LLMs) and Large Multimodal Models (LMMs), predominantly reason through textual representations, limiting their effectiveness in complex multi-step combinatorial and planning tasks. In this paper, we propose a zero-shot fully automatic framework which enables LMMs to reason through multiple chains of (self-generated) intermediate conceptual diagrams, significantly enhancing their combinatorial planning capabilities. Our approach does not require any human initialization beyond natural language description of the task. It integrates both textual and diagrammatic reasoning within an optimized graph-of-thought inference framework, enhanced by beam search and depth-wise backtracking. Evaluated on multiple challenging PDDL planning domains, our method substantially improves GPT-4o’s performance (eg., $35.5\%\rightarrow 90.2\%$ in Blocksworld). On more difficult planning domains with solution depths up to 40, our approach outperforms even the o1-preview reasoning model (eg., over 13% improvement in Parking). These results highlight the value of conceptual diagrams as a complementary reasoning medium in LMMs.
11footnotetext: Xero Inc.22footnotetext: BAIR, UC Berkeley33footnotetext: MIT-IBM Watson AI Lab
Figure 1: Our proposed approach. (1) Model generates an initial diagram schema and conceptual diagram for the domain, visually illustrating spatial information using simple shapes and colors. (2) Inference structured as a graph-of-thought with textual and diagrammatic representations for each state. (3) Model creates diagrams for initial and goal states of the instance conditioned on the conceptual diagram from step (1). (4) Generate each child state and its diagram based on its textual description and parent state. (5) Inference optimizations: Backtracking—if all candidate states at a depth fail verification, backtrack to deepest available ancestor; Beam search—rank candidate states by proximity to goal, selecting top $k=4$ for expansion (Sec. 2). (6) Stop when goal state is reached or maximum allowed states are visited. Example diagrams are from the Tetris domain, which requires rearranging tiles on a grid.
@@ 1 Introduction
Natural language is a powerful medium for communication, enabling humans to effectively share knowledge and ideas [44, 7, 49]. However, language alone is not an optimal medium for reasoning, as it is inherently linear, sequential, and verbose, making it inefficient for representing complex logical and relational structures [30, 22, 50]. Prior evidence [28, 16, 22, 4] indicates that humans do not rely solely on language when reasoning. Instead, they construct and utilize internal mental models—simplified analogues of real or hypothetical situations. Humans dynamically manipulate these mental models through mental simulations, which involve representing and predicting interactions between objects, to solve problems. Crucially, mental models are multimodal, integrating both visual and verbal representations to facilitate learning and robust reasoning [37]. Cognitive science research demonstrates that ‘human thought’ is not inherently verbal, sequential, or linear; rather, it is spatial, parallel, and image-like [50]. Finally, visual representations have always played a central role in human reasoning and communication, from prehistoric cave art—one of the earliest known forms of communication, predating written language [8]—to contemporary textbook diagrams, scientific figures in articles, and blackboard sketches.
@@
Sketches, better known as ‘conceptual diagrams’ in the cognitive science literature [30, 50, 22], externalize humans’ internal mental models, drawn to reduce cognitive load and facilitate rapid perceptual inferences and clearer reasoning. Instead of capturing fine-grained details of the external world, similar to the way human eyes perceive the environment, conceptual diagrams abstract away irrelevant details, use simple shapes and colors to represent objects, and emphasize intuitive spatial and relational interactions among entities. [50, 30]. For example, a square in a diagram can represent a complex real-world objects like a car, where the size, color, and position of the shape visually encode its relationships with other entities rather than the entity’s detailed visual appearance. Thus, conceptual diagrams can be an effective reasoning medium complementary to natural language, addressing language’s inherent limitations in accurate representation of spatial and relational structures and aligning closely with humans’ multimodal reasoning [50, 22].
@@
Modern large language models (LLMs) and large multimodel models (LMMs), such as GPT-4 [40] GPT4-o [39] and Gemini [19], have achieved remarkable success on complex mathematical and scientific benchmarks, including GSM8K [9], MATH [24], and GPQA [62]. Despite these advancements, their reasoning performance remains inconsistent, particularly on multi-step compositional reasoning, long-horizon planning, and tasks requiring backtracking and error correction [15, 53, 12]. These limitations partly arise because language models rely solely on language for reasoning, which is inherently inefficient for representing complex relational structures [15, 53]. Additionally, the autoregressive architecture of current LLMs enforces sequential next-token prediction, making tasks that require backtracking particularly challenging [15]. Thus, enabling LLMs to generate and reason with conceptual diagrams within graph-based inference frameworks that enable backtracking is an intuitive and promising approach to alleviate these bottlenecks.
@@
In this work, we propose a framework that enables LMMs to solve combinatorial planning problems by combining textual reasoning with self-generated conceptual diagrams. Our approach operates in a true zero-shot setting, requiring no domain-specific modifications or manual engineering to solve any combinatorial problem that can be expressed in Planning Domain Definition Language (PDDL) format [38], given only the natural language description of the problem (i.e., a clear textual specification of the initial state, goal state, and possible actions). Importantly, our method does not require human specification of the type or style of conceptual diagram to use, and is training-free. Our method works by decomposing inference into a structured graph of intermediate reasoning steps, where at each node the model determines the next best action to take from the current state. Specifically, each node in the inference graph is multimodal, containing both the textual description of the current state and a corresponding conceptual diagram visualizing that state. At each step of inference, the LMM: (i) generates the candidate next state; (ii) verifies uniqueness of the action chosen; (iii) produces a diagram schema — a structured set of textual statements specifying each object’s shape, size, location, and status in the diagram; (iv) Our model autonomously generates executable Matplotlib code for the conceptual diagrams, conditioned on the diagram schema. To ensure inference quality without sacrificing generality, we incorporate guardrails such as diagram schema self-reflection consistency checks and local (parent-child) and global validity checks (entire path from the initial state).
@@
Figure 2: We generate an initial conceptual diagram for each domain by sampling multiple diagram codes. An LMM ranks diagrams based on intuitive and accurate visualization of relational and spatial information. The top-ranked diagram’s code serves as reference for generating initial and goal state diagrams for all instances. Example shown is the parking domain, where curbs hold up to two cars, and cars can be movable or blocked (double-parked).
@@
Unlike other approaches that have explored augmenting language models with visual representations for compositional reasoning [59, 56, 26], our method generates conceptual diagrams entirely from scratch, without relying on human-provided initial visual demonstration of the domain or instance, and without requiring tuning of model’s visual encodings. Furthermore, rather than producing a single visualization, our approach autonomously generates multiple sequences of intermediate conceptual diagrams, explicitly illustrating how the LMM’s ‘model’ of the problem evolves with each reasoning step. To address the exponential growth of the combinatorial search space [5, 61], we employ beam search, where validated candidate states at each depth pf the inference graph are ranked by the model based on their proximity to the goal state, and only the top $k$ states are expanded further. Additionally, we incorporate a backtracking mechanism to avoid incomplete searches, allowing the model to revisit previously validated nodes if all candidate states at a given depth fail verification.
Evaluations across multiple challenging PDDL planning domains demonstrate that our method substantially enhances GPT-4o’s combinatorial reasoning capabilities. On the widely studied Blocksworld domain, from Valmeekam et al. [52], our approach significantly improves GPT-4o’s accuracy from 35.5% to 90.2%. Importantly, we contribute a new planning benchmark on five additional complex planning domains — Floor Tiles, Parking, Tetris, Elevator, and Barman —with solution depths up to 40, our method consistently and significantly outperforms GPT-4o, which typically fails entirely (e.g., 0% vs. 36.7% in Floor Tiles, 0% vs. 40% in Tetris). Furthermore, our approach consistently surpasses the best-performing reasoning model, o1-preview [42], across all evaluated complex domains (e.g., 26.7% vs. 40% in Tetris), highlighting the effectiveness of conceptual diagrams as a complementary reasoning medium, enabling LMMs to reliably solve challenging combinatorial planning problems that are otherwise beyond their capabilities, without requiring additional training or domain-specific engineering.
Figure 4: Sequence of intermediate state diagrams in the correct chain, from the initial state to the found goal state, for an example instance of five evaluated domains: Blocksworld, Floor Tile, Elevator, Parking, and Tetris.
2 Method
We propose a zero-shot framework (starting from text-only problem instance description) enabling Large Multi-Modal Models (LMMs), eg., GPT-4o, to solve combinatorial planning problems. Our framework integrates textual reasoning with (self-created by the model) intermediate visual diagrams, built upon a structured graph-of-thought inference approach. Combinatorial planning problems   involve finding a valid sequence of actions from an initial state $s_{0}$ to a goal state $s_{g}$, given a finite set of possible actions. Below, we detail our method following the stages illustrated in Fig. 1: starting from a textual description of the problem instance, (1) generating an initial domain-specific conceptual diagram, (2) visual graph-of-thought inference, (3) initial and goal state diagram generation (4) intermediate state generation pipeline, (5) inference optimizations (beam search and backtracking), and (6) search termination. Detailed prompts for all the steps are provided inside our code that is included an supplementary and would be released upon acceptance.
@@ 2 Method
See step 1 in Fig. 1. For each new planning domain, the LMM generates a domain-specific conceptual diagram that visually represents objects and their interactions using simple shapes, colors, and spatial arrangements. This is done entirely in a zero-shot manner, without manual engineering or external visual cues beyond the problem’s textual description. We begin by selecting a random instance from the domain and prompting the LMM to generate multiple candidate diagram schemas. The LMM then ranks these schemas based on how clearly they represent object positions and relationships, selecting the top-ranked one. Using this schema and the problem description, we generate multiple executable Matplotlib diagram codes to visualize the domain (Fig. 2). The LMM then ranks the rendered diagrams based on how effectively they visualize the spatial structure and relationships between objects. The highest-ranked diagram is chosen as the ‘example’ domain-specific conceptual diagram, and its code serves as a reference for generating the diagram code for the initial and goal states of all instances in the domain.
2. Visual Graph-of-Thought Inference.
See step 2 in Fig. 1. Our inference procedure decomposes reasoning into discrete steps, each represented as a node in a structured inference graph, inspired by Graph-of-Thought (GoT)  [5]. With slight abuse of notation we denote by $s_{d}$ an intermediate state at depth $d$ of the graph, represented by a combination of: (i) textual description $T(s_{d})$; (ii) a diagram $D(s_{d})$; and (iii) the action path $A_{0:d}$ from the initial state $s_{0}$. We iteratively expand the inference graph in a breadth-first search (BFS)  [10] manner, expanding nodes on the same depth following the beam search strategy explained in (5) below. Through this inference graph, we simultaneously explore multiple chains of intermediate states toward the goal state.
@@
See step 3 in Fig. 1. We begin the inference process by generating diagrams for the initial state $s_{0}$ and the goal state $s_{g}$ of each instance, conditioned on the domain conceptual diagram code obtained in step 1. The LMM is instructed to strictly adhere to the representation of objects and their statuses as illustrated in the domain’s conceptual diagram, while accurately initializing the objects according to the specific instance in the diagram code.
@@
See step 4 in Fig. 1. From each parent state $s_{d}$, we sample $n=4$ child states. W.l.o.g., next we describe the generation of a single such child state $s_{d+1}$ (Fig. 3). At each node, the LMM first selects the next candidate action $a_{d+1}$, conditioned on the parent state $s_{d}$ (which is represented by its diagram, textual description, and action path from initial state). We then generate the textual description $T(s_{d+1})$ of the resulting child state. The candidate action-state pair $(a_{d+1},T(s_{d+1}))$ is compared to previously generated child states to verify uniqueness. If different, the LMM generates a $s_{d+1}$ specific instance of the problem’s conceptual mapping $E(s_{d+1})$, which is a set of textual statements specifying shapes, sizes, positions, statuses (e.g., colors), and textual identifiers for each object in the state (see Fig. 3 for an example). The $E(s_{d+1})$ undergoes a self-reflection verification step to ensure consistency with $T(s_{d+1})$ and accuracy with respect to $a_{d+1}$, if failed we regenerate it. Subsequently, Matplotlib code $C(s_{d+1})$ is generated conditioned on $E(s_{d+1})$, $T(s_{d+1})$, and two example diagram codes: the initial state diagram code $C(s_{0})$ and the parent state diagram code $C(s_{d})$. Code is re-generated if it fails to run. After successfully generating the child state diagram, we perform two additional action verification checks: (1) a local check that verifies the validity of the action $a_{d+1}$ based solely on the diagrams and textual descriptions of the parent and child states (ensuring the action complies with domain constraints); and (2) a global check that verifies whether the entire action path $A_{0:d+1}$ remains feasible for reaching the goal state $s_{g}$. If either of these checks fails, the child state is marked as invalid.
5. Inference Optimizations: Beam Search and Backtracking.
@@
See steps 5 in Fig. 1. Applying Graph-of-Thought [5] to combinatorial planning problems via naively expanding all nodes (eg., standard BFS) would result in exponential growth of the search tree [46]. To mitigate this, we employ two inference optimizations: beam search and depth-wise backtracking. Using a method inspired by beam search  [35], we first expand all states at each depth $d$ using BFS, generating up to $N$ child states per state. The LMM then ranks all candidate child states at depth $d+1$ based on their proximity to the goal state $s_{g}$, selecting only the top $k=4$ states for further expansion. This depth-wise ranking solves the exponential growth problem. Moreover, this ranking system resembles human problem-solving strategies, where shallow state-specific heuristics are employed to estimate how close intermediate states are to the goal  [18, 6, 45]. Additionally, we implement a depth-wise backtracking mechanism. If all candidate child states at a given depth fail verification, we backtrack to the deepest available ancestor nodes at depth $d_{\text{max}}$ and attempt new expansions. We allow up to $B=2$ backtracking attempts to any given depth. If all $B$ attempts at depth $d_{\text{max}}$ fail, we mark nodes at that depth as invalid and backtrack further to the next deepest available nodes at depth $d_{\text{max2}}$, where $d_{\text{max2}}<d_{\text{max}}$.

### v2

# Visualizing Thought: Conceptual Diagrams Enable Robust Combinatorial Planning in LMMs

@@ Abstract
Human reasoning relies on constructing and manipulating mental models—simplified internal representations of situations that we use to understand and solve problems. Conceptual diagrams (e.g., a sketch drawn by a human to aid reasoning) externalize these mental models, abstracting irrelevant details to efficiently capture how entities interact with each other. In contrast, Large Language Models (LLMs) and Large MultiModal Models (LMMs) predominantly reason through text, limiting their effectiveness in complex multi-step tasks. In this paper, we propose Visual Thinking, a zero-shot framework that enables LMMs to reason through multiple chains of (self-generated) conceptual diagrams, significantly enhancing their combinatorial planning capabilities. Our approach does not require any human initialization beyond the natural language description of the task. It integrates both textual and diagrammatic reasoning within an optimized Graph-of-Thought inference framework, enhanced by beam search and depth-wise backtracking. Evaluated on multiple challenging PDDL planning domains, our method substantially improves LMMs’ performance (e.g., GPT-4o: $35.5\%\rightarrow 90.2\%$ in Blocksworld) and consistently outperforms other text-only search-based inference methods. On more difficult planning domains with solution depths up to 40, our approach outperforms even the o1-preview reasoning model (e.g., 16 percentage points improvement in Floor Tiles). These results highlight the value of conceptual diagrams as a reasoning medium in LMMs.
00footnotetext: 1Xero Inc.  2BAIR, UC Berkeley  3MIT-IBM Watson AI Lab
@@ 1 Introduction
Natural language is a powerful medium for communication, enabling humans to effectively share knowledge and ideas Pinker and Bloom (1990); Clark (1996); Tomasello (2010). However, language alone is not an optimal medium for reasoning, as it is inherently linear, sequential, and verbose, making it inefficient for representing complex logical and relational structures Larkin and Simon (1987); Hegarty (2004); Tversky (2011). Prior evidence Johnson-Laird (1983); Gentner and Stevens (2001); Hegarty (2004); Battaglia et al. (2013) indicates that humans do not rely solely on language when reasoning. Instead, they construct and utilize internal mental models—simplified analogues of real or hypothetical situations. Humans dynamically manipulate these mental models through mental simulations, which represent and predict interactions between objects, to solve problems. Crucially, mental models are multimodal, integrating both visual and verbal representations to facilitate learning and robust reasoning Mayer (2002). Cognitive science research demonstrates that ‘human thought’ is not inherently verbal, sequential, or linear; rather, it is spatial, parallel, and image-like Tversky (2011). Finally, visual representations have always played a central role in human reasoning and communication, from prehistoric cave art—one of the earliest known forms of communication, predating written language Clottes (2008)—to contemporary textbook diagrams, scientific figures in articles, and blackboard sketches.
@@
Conceptual diagrams are simplified visual representations that use basic shapes to capture how entities interact with each other while abstracting away irrelevant details Tversky (2011); Larkin and Simon (1987). These diagrams externalize the internal mental models of humans, reducing cognitive load and allowing rapid perceptual inference and clearer reasoning Hegarty (2004); Gentner and Stevens (2001). Unlike images, which capture the fine-grained details of how objects appear in the external world as perceived by the human eye, conceptual diagrams emphasize the structural and relational information needed for reasoning: they use simple shapes, colors, and relative positions and sizes to represent interactions among entities Tversky (2011); Larkin and Simon (1987). For example, a square in a diagram can represent a complex real-world object such as a car, where its size, color, and position visually encode its relationships with other entities abstracting away the car’s detailed appearance. Thus, conceptual diagrams are an effective reasoning medium complementary to natural language, addressing language’s inherent limitations in accurate representation of relational structures and aligning closely with humans’ multimodal reasoning Tversky (2011); Hegarty (2004); Dziri et al. (2023).
@@
Modern large language models (LLMs) and large multimodal models (LMMs), such as GPT-4 OpenAI (2023) GPT-4o OpenAI (2024a) and Gemini Google AI (2025), have achieved remarkable success on complex mathematical and scientific benchmarks, including GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), and GPQA Zhong et al. (2023). Despite these advancements, their reasoning performance remains inconsistent, particularly on multi-step compositional reasoning, long-horizon planning, and tasks requiring backtracking and error correction Dziri et al. (2023); Valmeekam et al. (2023b); Creswell et al. (2023). These limitations partly arise because language models rely solely on language for reasoning, which is inherently inefficient for representing complex relational structures Dziri et al. (2023); Valmeekam et al. (2023b). Additionally, the autoregressive architecture of current LLMs enforces sequential next-token prediction, making tasks that require backtracking particularly challenging Dziri et al. (2023). Thus, enabling LLMs to generate and reason with conceptual diagrams within graph-based inference frameworks that enable backtracking is an intuitive and promising approach to alleviate these bottlenecks.
Figure 1: Our proposed approach. Example diagrams are from the Tetris domain, where tiles are moved on a grid to reach a goal state. (1) The model generates multiple diagram schemas and codes for a random instance; their rendered diagrams are ranked; the code of the top choice is cached (Fig. 2). (2) Conditioned on this code, the model produces diagrams for the initial and goal states. (3) Candidate child state, along with its diagrams, is generated based on the parent state (Fig. 3). (4) Beam search ranks all candidates at a depth by proximity to the goal, expanding the top $k=4$ (Sec. 3). (5) Depth-wise backtracking is applied when all candidate states at a depth fail validation, returning to the deepest available ancestor. (6) The process stops when the goal is reached or a maximum number of steps is visited. (7) The output is the action sequence (plan) plus textual and diagrammatic representations of intermediate states.1
11footnotetext: For Figs. 1 and 4, we adjusted the diagrams to increase font sizes for better readability. Figs. 2-3 are unmodified generations. Fig. 1 action descriptions are simplified; full action strings include previous/current cells occupied.
@@
In this work, we propose a framework that enables LMMs to solve combinatorial problems by combining textual reasoning with self-generated conceptual diagrams. Our approach operates in a zero-shot setting, requiring no domain-specific modifications or manual engineering to solve any combinatorial problem that can be expressed in Planning Domain Definition Language (PDDL) format McDermott et al. (1998), given only the natural language description of the problem (i.e., a textual specification of the problem’s initial state, goal state, and possible actions). Importantly, our method does not require any visual templates or geometric specifications; instead, it generates conceptual diagrams entirely from the textual description of the problem.
Our method decomposes inference into a graph of intermediate reasoning steps, where at each node the model determines the next best state from the current state. Each node is multimodal, containing both the textual description of the state and a corresponding conceptual diagram. At each step, the LMM: (i) generates the candidate next state using the textual and diagrammatic representations of previous states; (ii) produces a diagram schema—a structured set of textual statements specifying each object’s shape, relative size and location, and status; and (iii) autonomously generates executable Matplotlib code that renders the conceptual diagram of the state, conditioned on the schema. To ensure inference quality, we incorporate guardrails such as diagram schema self-reflection checks and local (parent-child) and global (entire path from the initial state) validity checks.
Unlike prior approaches that augment language models with visual representations for compositional reasoning Wu et al. (2024); Wang et al. (2024); Hu et al. (2024)—where models are given an initial visual demonstration or template of the state and then asked to update it—our method relies solely on the textual description of each state. The model independently generates conceptual diagrams for every state entirely from scratch, without any human-provided visual examples, templates, or cues, and without any tuning of the model’s visual encodings. This end-to-end diagram generation process mirrors how humans use imagination to construct mental models from language. Furthermore, rather than producing a single visualization, our approach autonomously generates multiple sequences of intermediate conceptual diagrams, explicitly illustrating how the LMM’s ‘model’ of the problem evolves with each reasoning step. To address the exponential growth of the combinatorial search space Besta et al. (2023); Yao et al. (2023), we employ beam search, where validated candidate states at each depth of the inference graph are ranked by the model based on their proximity to the goal state, and only the top $k$ states are expanded further. Additionally, we incorporate a backtracking mechanism to avoid incomplete searches, allowing the model to revisit previously validated nodes if all candidate states at a given depth fail verification.
Evaluations across multiple challenging PDDL planning domains demonstrate that our method substantially enhances LLMs’ combinatorial reasoning capabilities. On the widely studied Blocksworld domain Valmeekam et al. (2023a), our approach yields performance gains of roughly 43, 64, and 55 percentage points using Claude 3.5 Sonnet Anthropic (2024), LLaMA 4 Maverick Meta (2025), and GPT-4o, respectively (e.g., GPT-4o’s accuracy rises from 35.5% to 90.2%). Importantly, we contribute a new planning benchmark on five additional domains—Floor Tiles, Parking, Tetris, Elevator, and Barman—with solution depths designed up to 40, where our method consistently succeeds even as the base model (GPT-4o) fails entirely (e.g., 0% vs. 36% in Floor Tiles, 0% vs. 38% in Tetris). Furthermore, our approach with GPT-4o as the base model, surpasses the best-performing reasoning model, o1-preview OpenAI (2024c), across all evaluated complex domains (e.g., 26% vs. 38% in Tetris). Finally, compared to strong search-based planning methods such as Graph-of-Thought and RAP (a Monte Carlo Tree Search framework that uses an LLM both as world model and planner), our method improves accuracy by at least 22 percentage points while also reducing inference cost by over 30% and latency by more than 25%. These results highlight conceptual diagrams as an efficient reasoning medium, enabling LMMs to reliably solve challenging long-horizon planning problems that remain beyond the reach of purely textual (single inference or search-based) approaches.
Single inference Search-based inference (base LLM = GPT-4o)
Domain
GPT-4o
o1-
mini
o1-
preview
GoT
Optimized
GoT
RAP(20)
Visual Thinking
Blocksworld (simple) 35.5% 56.6% 97.8% 50% 58% 58% 90.2%
Blocksworld (hard) 0% 0.9% 23.65% 8% 48% 4% 78%
Floor Tiles 0% 6% 20% 0% 4% 36%
Parking 2% 8% 40% 14% 28% 52%
Tetris 0% 2% 26% 0% 12% 38%
Elevator 0% 2% 36% 2% 10% 48%
Barman 0% 0% 10% 0% 4% 30%
Table 1: Accuracy results across all evaluated domains. Each baseline was evaluated on 50 problem instances per domain, except for GPT-4o + Visual Thinking on Blocksworld (simple), where we evaluated 500 instances (the full PlanBench size Valmeekam et al. (2023a)) and achieved 90.2% (451/500). Baseline results marked with are taken from Valmeekam et al. (2024). RAP reported 51% accuracy on Blocksworld (simple) using LLaMA 2 Hao et al. (2023).
2 Related Work
Multimodal Chain-of-Thought for Reasoning. Several recent works have explored integrating visual representations into reasoning processes for LLMs and LMMs. Hu et al. (2024) equips multimodal models with drawing tools to generate intermediate sketches composed of lines and shapes, but primarily focuses on single-step or shallow problems, typically producing one intermediate visualization per instance. Similarly, Wang et al. (2024) generates visual aids for spatial reasoning, providing a single refined visualization per problem. In both works, generated visualizations are typically an approximate or an augmentation of a high-fidelity illustration rather than conceptual diagrams representing object interactions. Concurrent work, Wu et al. (2024), generates visual and textual intermediate steps for planning tasks by conditioning multimodal models on externally provided initial visual representations that supplement textual problem descriptions. Their approach iteratively generates intermediate textual and visual states, but the reliance on externally supplied initial visualizations could limit generality and applicability to unseen domains.
Our approach differs in several key aspects. First, our method autonomously generates conceptual diagrams directly from textual problem descriptions, without relying on external visual demonstrations or cues, mirroring the human ability to construct mental models from language. Second, rather than producing a single visualization, our approach creates multiple chains of intermediate visual states, enabling parallel multi-hypothesis compositional reasoning through evolving diagrams. Third, our diagrams explicitly represent relationships between objects in a conceptual manner, rather than realistic or approximate visual depictions of the objects. Finally, our method operates in a zero-shot setting, and is applicable to any problem expressible in PDDL format (a general language for planning problems) without any domain-specific information required. These distinctions collectively enable a more generalizable and flexible form of diagrammatic reasoning, leading to significant performance gains.
Search and Verification Strategies for Enhanced LLM Reasoning. Recent methods improve reasoning in LLMs by structuring inference into explicit intermediate steps Wei et al. (2022); Kojima et al. (2022) and employing search-based strategies over multiple reasoning paths Wang et al. (2022); Yao et al. (2023); Hao et al. (2023); Creswell et al. (2022); Besta et al. (2023). For instance, Yao et al. (2023) propose Tree-of-Thought (ToT), which extends CoT to explore a tree of reasoning paths, and Besta et al. (2023) introduce Graph-of-Thought (GoT), which structures the reasoning process as a graph, enabling exploration and aggregation of intermediate reasoning steps. Other works utilize verification and refinement techniques, such as iterative self-reflection and feedback, to enhance reasoning accuracy Madaan et al. (2023); Paul et al. (2023); Shinn et al. (2023). Our approach builds upon these methods by integrating conceptual diagrams into intermediate reasoning steps. Moreover, as detailed in Section 3, we extend graph-based inference with beam search and depth-wise backtracking, significantly reducing the search space and improving performance on long-horizon planning tasks.
World Modeling and Planning with Language Models. Recent research explores planning and reasoning using LLMs by implicitly or explicitly constructing world models from textual descriptions Huang et al. (2022); Yao et al. (2022); Ahn et al. (2022); Dasgupta et al. (2022); Driess et al. (2023); Liu et al. (2023a); Wang et al. (2023). For instance, Huang et al. (2022) show that LLMs implicitly form textual world models for simple planning tasks, while other works explicitly represent states and transitions for structured reasoning Yao et al. (2022); Dasgupta et al. (2022); Liu et al. (2023a). Recent approaches also introduce explicit state-tracking, modeling entity relationships and transitions to enhance reasoning Liu et al. (2023a); Wang et al. (2023); Driess et al. (2023). Our work extends these approaches by enabling LMMs to autonomously construct and reason with diagrammatic world models. Specifically, our method generates conceptual diagrams directly from textual problem descriptions, defining a visual schema of objects and their relationships through abstract shapes, colors, and spatial relationships. Subsequent diagram updates visually simulate action sequences, explicitly tracking object states and interactions. Compared to purely textual world modeling, our diagrammatic approach provides clearer and more accurate representations of spatial and relational information, significantly improving performance on complex, multi-step planning tasks.
@@ 2 Related Work
Figure 2: We generate an initial conceptual diagram for each domain by sampling multiple diagram codes. An LMM ranks diagrams based on intuitive and accurate visualization of relational information. The top-ranked diagram’s code serves as reference for generating initial and goal state diagrams for all instances. Example shown is the parking domain, where curbs hold up to two cars, and cars can be movable or blocked (double-parked).
3 Method
We propose Visual Thinking, a training-free and model-agnostic framework, which integrates textual reasoning with model-generated intermediate conceptual diagrams, to enable Large MultiModal Models (LMMs) to solve combinatorial problems, given text-only problem specifications (initial state, goal, admissible actions). Combinatorial problems  Wikipedia (2025) involve finding a valid sequence of actions from an initial state $s_{0}$ to a goal state $s_{g}$, given a finite set of possible actions. Our framework, built upon the Graph-of-Thought inference approach  Besta et al. (2023), decomposes reasoning into discrete steps, each represented as a node in a structured inference graph. Through this inference graph, we simultaneously explore multiple chains of intermediate multimodal states toward the goal state. Below, we detail our method following the stages illustrated in Fig. 1. The full implementation, including prompts and outputs, is available in the paper’s GitHub repository.
@@ 3 Method
See step 1 in Fig. 1. For each new planning domain, the LMM generates a domain-specific conceptual diagram that visually represents objects and their interactions using simple shapes, colors, and spatial arrangements. This is done entirely in a zero-shot manner, without manual engineering or external visual cues, conditioned only on the problem’s textual description. We begin by selecting a random instance from the domain and prompting the LMM to generate multiple candidate diagram schemas. The LMM then ranks these schemas based on how clearly they represent object positions and relationships, selecting the top-ranked one. Using this schema and the problem description, we generate multiple executable Matplotlib diagram codes to visualize the domain (Fig. 2). The LMM then ranks the rendered diagrams based on how effectively they visualize the structure and relationships between objects. The highest-ranked diagram is chosen as the ‘example’ domain-specific conceptual diagram, and its code serves as a reference for generating the diagram code for the initial and goal states of all instances in the domain.
@@
See step 2 in Fig. 1. We begin the inference process by generating diagrams for the initial state $s_{0}$ and the goal state $s_{g}$ of each instance, conditioned on the domain conceptual diagram code obtained in step 1. The LMM is instructed to strictly adhere to the representation of objects and their statuses as illustrated in the domain’s conceptual diagram, while accurately initializing the objects according to the specific instance in the diagram code.
@@
See step 3 in Fig. 1. We denote by $s_{d}$ an intermediate state at depth $d$ of the graph, represented by a combination of: (i) textual description $T(s_{d})$; (ii) a diagram $D(s_{d})$; and (iii) the action path $A_{0:d}$ from the initial state $s_{0}$. We iteratively expand the inference graph depth by depth in a breadth-first search (BFS)  Cormen et al. (2009) manner, and apply beam search at each depth to select the top-k candidates for further expansion.
@@
From each parent state $s_{d}$, we sample $n=4$ child states. W.l.o.g., next we describe the generation of a single such child state $s_{d+1}$ (Fig. 3). At each node, the LMM first selects the next candidate action $a_{d+1}$, conditioned on the parent state $s_{d}$ (which is represented by its diagram, textual description, and action path from initial state). We then generate the textual description $T(s_{d+1})$ of the resulting child state. The candidate action-state pair $(a_{d+1},T(s_{d+1}))$ is compared to previously generated child states to verify uniqueness. If different, the LMM generates a $s_{d+1}$ specific instance of the problem’s conceptual mapping $E(s_{d+1})$, which is a set of textual statements specifying shapes, sizes, positions, statuses (e.g., colors), and textual identifiers for each object in the state (see Fig. 3 for an example). The $E(s_{d+1})$ undergoes a self-reflection verification step to ensure consistency with $T(s_{d+1})$ and accuracy with respect to $a_{d+1}$, if failed we regenerate it. Subsequently, Matplotlib code $C(s_{d+1})$ is generated conditioned on $E(s_{d+1})$, $T(s_{d+1})$, and two example diagram codes: the initial state diagram code $C(s_{0})$ and the parent state diagram code $C(s_{d})$. Code is re-generated if it fails to run. After generating the child state diagram, we perform two additional action verification checks: (1) a local check that verifies the validity of the action $a_{d+1}$ based solely on the diagrams and textual descriptions of the parent and child states (ensuring the action complies with domain constraints); and (2) a global check that verifies whether the entire action path $A_{0:d+1}$ remains feasible for reaching the goal state $s_{g}$. If either of these checks fails, the child state is marked as invalid.
4. Inference Optimization: Beam Search.
@@
See step 4 in Fig. 1. Applying Graph-of-Thought Besta et al. (2023) to combinatorial planning problems via naively expanding all nodes (e.g., BFS) would result in exponential growth of the search tree Russell and Norvig (2010). To mitigate this, we use a method inspired by beam search  Lowerre (1976), where first all states are expanded at each depth $d$ using BFS, generating up to $N$ child states per state. The LMM then ranks all candidate child states at depth $d+1$ based on their proximity to the goal state $s_{g}$, selecting only the top $k=4$ states for further expansion. This depth-wise ranking solves the exponential growth problem. Moreover, this ranking system resembles human problem-solving strategies, where shallow state-specific heuristics are employed to estimate how close intermediate states are to the goal  Gigerenzer and Goldstein (1996); Chronicle et al. (2006); Pizlo and Li (2005).
Figure 4: Sequence of intermediate state diagrams in the correct chain, from the initial state to the found goal state, for an example instance of five evaluated domains: Blocksworld, Floor Tiles, Elevator, Parking, and Tetris.
5. Depth-wise Backtracking.
See step 5 in Fig. 1. Additionally, we implement a depth-wise backtracking mechanism. If all candidate child states at a given depth fail verification, we backtrack to the deepest available ancestor nodes at depth $d_{\text{max}}$ and attempt new expansions. We allow up to $B=2$ backtracking attempts to any given depth. If all $B$ attempts at depth $d_{\text{max}}$ fail, we mark nodes at that depth as invalid and backtrack further to the next deepest available nodes at depth $d_{\text{max2}}$, where $d_{\text{max2}}<d_{\text{max}}$.